ANALYSIS OF HIGH-RESOLUTION SPECTRA OF CYCLOPROPANE

\underline{TONY MASIELLO}, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, Richland, WA 99354 (PNNL is operated for the US Department of Energy by the Battelle Memorial Institute under contract DE-AC05-76RLO 1830); MARK J. ABEL, JOSEPH W. NIBLER, Department of Chemistry, Oregon State University, Corvallis, OR 97331.

Cyclopropane is the simplest of the strained carbon compounds that form rings or cages. Though the lowest vibrational state lies at 739 cm^{−1}, half of the infrared active fundamental vibrations are perturbed by some form of a combination and/or overtone vibration. Previous work was restricted in that the majority of the lower vibrational states are either dark or infrared inactive. This forced investigators into giving their best estimates of the rotational constants of perturbing states and in some cases their best estimates as to the identification of the perturbing states. This has prompted a reinvestigation of cyclopropane using high-resolution infrared spectroscopy. A Coriolis interaction that couples ν_{7} to ν_{11} was identified and an improved set of rovibrational constants will be given. In addition, a high-resolution CARS spectrum of the ν_{2} Raman band that is involved in a Fermi resonance triad with 2ν_{14} and ν_{9} provides an additional check to the validity of the assignments.